# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v2

@@ Appendix C Equivariant Brouwer Degree Background
Equivariant notation. Let $\mathcal{G}$ be a compact Lie Group. For any subgroup $H\leq\mathcal{G}$, denote by $(H)$ its conjugacy class, by by $N(H)$ its normalizer by $W(H):=N(H)/H$ its Weyl group in $\mathcal{G}$. The set of all subgroup conjugacy classes in $\mathcal{G}$, denoted $\Phi(\mathcal{G}):=\{(H):H\leq\mathcal{G}\}$, has a natural partial order defined as follows
$(H)\leq(K)\iff\exists_{g\in\mathcal{G}}\;\;gHg^{-1}\leq K.$
In particular, we put $\Phi_{0}(\mathcal{G}):=\{(H)\in\Phi(\mathcal{G})\;:\;\text{$W(H)$ is finite}\}$. For a $\mathcal{G}$-space $X$ and $x\in X$, denote by $\mathcal{G}_{x}:=\{g\in\mathcal{G}:gx=x\}$ the isotropy group of $x$ and call $(\mathcal{G}_{x})\in\Phi(\mathcal{G})$ the orbit type of $x\in X$. Put $\Phi(\mathcal{G},X):=\{(H)\in\Phi_{0}(\mathcal{G})\;:\;(H)=(\mathcal{G}_{x})\;\text{for some $x\in X$}\}$ and $\Phi_{0}(\mathcal{G},X):=\Phi(\mathcal{G},X)\cap\Phi_{0}(\mathcal{G})$. For a subgroup $H\leq\mathcal{G}$, the subspace $X^{H}:=\{x\in X:\mathcal{G}_{x}\geq H\}$ is called the $H$-fixed-point subspace of $X$. If $Y$ is another $\mathcal{G}$-space, then a continuous map $f:X\to Y$ is said to be $\mathcal{G}$-equivariant if $f(gx)=gf(x)$ for each $x\in X$ and $g\in\mathcal{G}$. The Burnside Ring and Axioms of Equivariant Brouwer Degree. Denote by $\mathcal{M}^{G}$ the set of all admissible $G$-pairs by $A(G)$ the Burnside ring of $G$. The following statement is the standard axiomatic definition of the $G$-equivariant Brouwer degree (cf. [5]).

### v3

@@ Appendix C Equivariant Brouwer Degree Background
We encourage readers interested in exploring the statements made in this section with greater detail to refer to the texts [4], [5]. Equivariant notation. Let $G$ be a compact Lie Group. For any subgroup $H\leq G$, denote by $(H)$ its conjugacy class, by by $N(H)$ its normalizer by $W(H):=N(H)/H$ its Weyl group in $G$. The set of all subgroup conjugacy classes in $G$, denoted $\Phi(G):=\{(H):H\leq G\}$, has a natural partial order defined as follows
$(H)\leq(K)\iff\exists_{g\in G}\;\;gHg^{-1}\leq K.$
In particular, we put $\Phi_{0}(G):=\{(H)\in\Phi(G)\;:\;\text{$W(H)$ is finite}\}$ and, for any $(H),(K)\in\Phi_{0}(G)$, we denote by $n(H,K)$ the number of subgroups $\tilde{K}\leq G$ with $\tilde{K}\in(K)$ and $H\leq\tilde{K}$. For a $G$-space $X$ and $x\in X$, denote by $G_{x}:=\{g\in G:gx=x\}$ the isotropy group of $x$ and call $(G_{x})\in\Phi(G)$ the orbit type of $x\in X$. Put $\Phi(G,X):=\{(H)\in\Phi_{0}(G)\;:\;(H)=(G_{x})\;\text{for some $x\in X$}\}$ and $\Phi_{0}(G,X):=\Phi(G,X)\cap\Phi_{0}(G)$. For a subgroup $H\leq G$, the subspace $X^{H}:=\{x\in X:G_{x}\geq H\}$ is called the $H$-fixed-point subspace of $X$. If $Y$ is another $G$-space, then a continuous map $f:X\to Y$ is said to be $G$-equivariant if $f(gx)=gf(x)$ for each $x\in X$ and $g\in G$. [5]. The Burnside Ring and Axioms of Equivariant Brouwer Degree. The free ${\mathbb{Z}}$-module $A(G):={\mathbb{Z}}[\Phi_{0}(G)]$ becomes the Burnside ring when equipped with a natural multiplicative operation, called the Burnside ring product and defined for any pair of generators $(H),(K)\in\Phi_{0}(G)$ as follows
(59) $(H)\cdot(K):=\sum\limits_{(L)\in\Phi_{0}(G)}n_{L}(L),$
where the coefficients $n_{L}\in{\mathbb{Z}}$ are given by the recurrence formula
(60) $n_{L}:=\frac{n(L,H)|W(H)|n(L,K)|W(K)|-\sum_{(\tilde{L})>(L)}n_{\tilde{L}}n(L,\tilde{L})|W(\tilde{L})|}{|W(L)|}.$
Let $V$ be an orthogonal $G$-representation with an open bounded $G$-invariant set $\Omega\subset V$. A $G$-equivariant map $f:V\rightarrow V$ is said to be $\Omega$-admissible if $f(x)\neq 0$ for all $x\in\partial\Omega$, in which case the pair $(f,\Omega)$ is called an admissible $G$-pair in $V$. We denote by $\mathcal{M}^{G}(V)$ the set of all admissible $G$-pairs in $V$ and by $\mathcal{M}^{G}$ the set of all admissible $G$-pairs defined by taking a union over all orthogonal $G$-representations, i.e.
$\mathcal{M}^{G}:=\bigcup\limits_{V}\mathcal{M}^{G}(V).$
The following statement is the standard axiomatic definition of the $G$-equivariant Brouwer degree: